5) Prove by induction. For every integer n 23, 5(5 - 25) 5° +54 + ........... +5 = 4
Answer #1

ie, we prove that 554...+ sh 5kt_56507 25) let sin) 53+54+---+S= s(s-25) 4 for na LHS = 53= 125 and R:HS = 5(8-25) 5 (125-25